OpenStudy (anonymous):

World problem Algebra 2. A mom buys 3 different types if t-shirts for her children, long sleeve, polo, and short sleeve. The long sleeve shirts cost $18 each, the polos cost $15 each, and the short sleeve cost $12 each. The mother purchased 3 less polos then she did short sleeve shirts. If the mother purchased 18 shirts at an average of $14 each, find the number of each type of shirt purchased

OpenStudy (anonymous):

@satellite73 could you help me me set of the equations?

OpenStudy (anonymous):

start by declaring your variable say "let \(x\) = ..."

OpenStudy (anonymous):

i set up p= s-3 and 18L + 15p+ 12s = 18 but Ihave a feelings there not right

OpenStudy (anonymous):

no it is not right \[18L+15P+12S\] represents the cost, not the total number of shirts bought

OpenStudy (anonymous):

if \(L\) is the number of long sleeve shirts, then the cost of all the long sleeve shirts is \(18L\) etc the total number of all the shirts is \[P+L+S\], which you know to be 18 so you can write \[P+L+S=18\] as one equation

OpenStudy (anonymous):

now we have to figure out how much was spent all together. she bought 18 shirts at an average of $14 per shirt so that shouldn't be too hard

OpenStudy (anonymous):

let me know when you get the answer (by multiplying)

OpenStudy (anonymous):

252

OpenStudy (anonymous):

ok good, so now we have 3.6*5 \[P+L+S=18\]\[18L+15P+12S=252\]and also \[P=S-3\]

OpenStudy (anonymous):

ignore that weird number up top, it is a type

OpenStudy (anonymous):

now lets go back and replace all the \(P\) what we see by \(S-3\)

OpenStudy (anonymous):

\[S-3+L+S=18\]\[15(S-3)+18L+12S=252\]

OpenStudy (anonymous):

we still have a bit more work to do because this problem is annoying first equation become \[2S+L-3=18\] or \[2S+L=21\]

OpenStudy (anonymous):

is that our final equation?

OpenStudy (anonymous):

oh no we also have the second equation which (after multiplying out and collecting terms) will be \[18L+27S=297\]

OpenStudy (anonymous):

whew this is a pain right? we have to solve the two equation \[2S+L=21\]\[18L+27S=297\] so we can do this by rewriting the first equation as \[L=21-2S\]substitute in the second one to get \[18(21-2S)+27S=297\] and FINALLY we have one equation to solve

OpenStudy (anonymous):

hopefully after a bit of algebra you will get \(S=9\)

OpenStudy (anonymous):

and of course we know S is short, so P is \(9-3=6\) and the rest are \(L\)